Consider the points below. 

P(1, 0, 1), ((-2, 1, 3), R(4, 2, 5)

 (a) Find a nonzero vector orthogonal to the plane through the points P, Q, and R.

 (b) Find the area of the triangle PQR. 
Answer #1

(START toly by p= (1,0,1), Q=(2,1,3), R = 14, 2,5) Pa = (2-1) + (1-0)j + (3-1) k = -sitj+2k a(-3,1,2) PR = (4-1); +(2-0)j +5-